OpenStudy (anonymous):

A vine 6cm. tall is planted in a garden. It grows 4cm. a day. How many days will it take to reach 112cm. in height?

OpenStudy (eli_moses):

for this problem you need an equation first

OpenStudy (eli_moses):

the starting equation for the plant is \[112=4x+6\] because you need to get to 112, the plant starts at 6 cm, and it rises 4 cm per day

OpenStudy (eli_moses):

now try to sole the equation

OpenStudy (anonymous):

x=53/2 ?

OpenStudy (eli_moses):



OpenStudy (anonymous):

26.5

OpenStudy (eli_moses):

yes so it will take 26.5 days

OpenStudy (anonymous):

thanks :)

OpenStudy (eli_moses):

anytime

OpenStudy (anonymous):

Which of the following is not a literal equation? a.x=y+3 b.x^2+y^2=z^2 c.n+n=10 d.(a/b)=1

OpenStudy (eli_moses):

which equation wouldn't be linear in this case

OpenStudy (eli_moses):

we see A is linear because it's equation is the same as y=x-3

OpenStudy (eli_moses):

which other one would be linear

OpenStudy (anonymous):

thats what we are asking you.

OpenStudy (eli_moses):

B has exponents so it's not linear on a graph the lines would be curved therefore, which is not a literal equation?

OpenStudy (anonymous):

Ok, thank you very much! I understand now!

OpenStudy (eli_moses):

no problem